# IndiSeek learns information-guided disentangled representations

### v1

@@ Abstract
Learning disentangled representations is a fundamental task in multi-modal learning. In modern applications such as single-cell multi-omics, both shared and modality-specific features are critical for characterizing cell states and supporting downstream analyses. Ideally, modality-specific features should be independent of shared ones while also capturing all complementary information within each modality. This tradeoff is naturally expressed through information-theoretic criteria, but mutual-information–based objectives are difficult to estimate reliably, and their variational surrogates often underperform in practice. In this paper, we introduce IndiSeek, a novel disentangled representation learning approach that addresses this challenge by combining an independence-enforcing objective with a computationally efficient reconstruction loss that bounds conditional mutual information. This formulation explicitly balances independence and completeness, enabling principled extraction of modality-specific features. We demonstrate the effectiveness of IndiSeekon synthetic simulations, a CITE-seq dataset and multiple real-world multi-modal benchmarks.
@@ 1 Introduction
The growing availability of multi-modal data has broadened the scope of representation learning. One notable example is the development of single-cell multi-omics technologies in genomics (Teichmann and Efremova,, 2020). For instance, CITE-seq (Cellular Indexing of Transcriptomes and Epitopes by Sequencing) (Stoeckius et al.,, 2017) enables simultaneous measurement of gene expression via single-cell RNA sequencing and protein abundance via antibody-derived tags (ADTs). Integrative analysis of both modalities at the single-cell resolution has improved tasks such as cell state delineation and enabled cell atlas construction and querying at finer granularity. Multi-modal data is also abundant at the intersection of computer vision, natural language processing, and audio processing, where combining information from different sources for representation learning forms the backbone of many groundbreaking progresses (Radford et al.,, 2021; Baltrušaitis et al.,, 2018; Jia et al.,, 2021; Akbari et al.,, 2021; Liang et al.,, 2024).
@@
Shared features enhance predictive power across modalities by capturing cross-modal dependence. From an information-theoretic perspective, they should contain all information common across modalities. zhe pu At the same time, preserving modality-specific diversity is equally important. In single-cell multi-omics, modality-specific features often correspond to unique biological signals that aid cell type identification and the discovery of novel or rare cell populations (Hao et al.,, 2021; Caron et al.,, 2025). After all, if all the information in a certain modality can be inferred from other modalities, there is little reason to measure it in the first place. Finally, disentangling shared and modality-specific components improves interpretability: shared features summarize cross-modal information, while modality-specific features capture complementary signals.
Methodologically, shared information is usually extracted by maximizing cross-modal mutual information (Tosh et al.,, 2021; Sridharan and Kakade,, 2008), with contrastive methods like CLIP (Radford et al.,, 2021) widely adopted in practice. Under certain conditions, CLIP-learned representations have been shown to be not only sufficient (i.e., capturing all shared information), but also minimally sufficient (i.e., containing minimal extra information) (Gui et al.,, 2025; Oko et al.,, 2025; Lin and Mei,, 2025) under certain conditions.
In contrast, identifying modality-specific features, also known as disentangled representation learning, has become an active research area (Fischer,, 2020; Liang et al.,, 2023; Liu et al.,, 2023; Wang et al., 2024a, ; Dufumier et al.,, 2024; Wang et al., 2024b, ). While shared features should be minimally sufficient, modality-specific features must both remain independent of (i.e., disentangled from) shared features and capture all complementary information within each modality (Liang et al.,, 2023; Wang et al., 2024a, ; Wang et al., 2024b, ). From an information-theoretic perspective, this requires (1) simultaneously minimizing the mutual information between shared and modality-specific features and (2) maximizing the mutual information between data and their union. Multiple state-of-the-art methods (Liang et al.,, 2023; Wang et al., 2024a, ) tackle this problem from such a perspective, but differ in how they approximate the underlying mutual information. Since such quantities are difficult to estimate with finite samples, these methods often trade off independence against completeness. As we show in the motivating examples below, this tradeoff limits their ability to achieve both goals simultaneously.
@@ 1.1 Motivating examples and limitations of SOTA
For any learned map $h_{1}$, we quantify the importance of each coordinate of $X_{1}$ in determining its output via gradients based on feature masking [^1] See Appendix A.1 for details on this feature importance metric. on an independent test set of size $1000$, following standard model-free importance measures (Robnik-Šikonja and Kononenko,, 2008; Zeiler and Fergus,, 2014; Li et al.,, 2016)
@@
We present the average coordinate importance metrics in learned modality-specific representation maps over $50$ simulation runs in Figures 1 and 2 for three different learning algorithms in the two settings, respectively. In both figures, the left panels report results of the SOTA approach FactorizedCL in Liang et al., (2023) (without self-supervision), the middle panels report results of the SOTA approach InfoDisen in Wang et al., 2024a , and the right panels report results of the new approach IndiSeek we propose in this paper. Tuning parameters [^2] See Appendix A.2 for an ablation study on performances with different tuning parameter values. are set at $0.01$ for InfoDisen and IndiSeek and at $1$ by default for FactorizedCL.
As we have reasoned when introducing the settings, the ideal modality-specific representations in Setting 1 should depend on and only on the last four coordinates, while in Setting 2, on and only the last two coordinates. The left and the middle panels of both figures suggest that neither FactorizedCL nor InfoDisen learns an ideal representation in either setting, as irrelevant coordinates play non-trivial roles in determining the learned modality-specific representations. Therefore, both SOTA approaches have their limitations in achieving the desired goal of disentangled representation learning. In contrast, the right panels of Figures 1 and 2 show that the new method IndiSeekwe are to propose achieves it in both settings: the learned modality-specific features depend on and only on all coordinates that are independent of those involved in shared features.
@@ 3 Information-guided disentangled representation seeking (IndiSeek)
Given learned shared features, the desiderata in Section 2 motivate the search for disentangled modality-specific features, but leave open the question of how to realize them. Prior work has proposed an information-theoretic objective (e.g., Wang et al., 2024a, ):
@@ 3.1 Step 1: Extracting Shared Features via CLIP
As noted at the end of Section 2, shared feature extraction is commonly handled by contrastive learning methods such as CLIP (Radford et al.,, 2021), which we also adopt. Concretely, encoders $(f_{1},f_{2})$ are trained with the InfoNCE loss (Oord et al.,, 2018), which provides a lower bound on the mutual information between $X_{1}$ and $X_{2}$. This encourages $C_{1}=f_{1}(X_{1})$ and $C_{2}=f_{2}(X_{2})$ to capture all cross-modal dependence while avoiding modality-specific redundancy.
@@ 3.2 Step 2: Extracting Modality-Specific Features
Given the shared features $(C_{1},C_{2})$, we aim to extract the modality-specific components $(Z_{1},Z_{2})$ by operationalizing the constrained problem in Eq. (3), which can be equivalently expressed in Lagrangian form as
$\min_{Z_{i}}\;\underbrace{I(Z_{i};C_{i})}_{\text{disentanglement-enforcing term}}-\quad\lambda\cdot\underbrace{I(Z_{i},C_{i};X_{i})}_{\text{complementary information capture term}}.$ (4)
@@
  • For the disentanglement enforcing term $I(Z_{i};C_{i})$, which we aim to minimize, we replace it by an upper bound given by the NCE-CLUB loss (Cheng et al.,, 2020; Liang et al.,, 2023):
    ${\mathcal{L}}_{\rm NCE-CLUB}(Z_{1};C_{1})=\mathbb{E}_{(Z_{1},C_{1})}\left[\log p(Z_{1}\mid C_{1})\right]-\mathbb{E}_{Z_{1},\widetilde{C}_{1}}\left[\log p(Z_{1}\mid\widetilde{C}_{1})\right],$ (5)
    where $p(z_{1}\mid c_{1})$ is the underlying conditional density of $Z_{1}\mid C_{1}$ and $\widetilde{C}_{1}$ is an independent copy of $C_{1}$ from the same marginal distribution. [^4] In implementation, following Liang et al., (2023), we replace conditional distributions by their estimates.
@@
$\mathcal{L}_{\text{{IndiSeek}}}(Z_{i};C_{i},\lambda)\;=\;\mathcal{L}_{\text{NCE-CLUB}}(Z_{i};C_{i})\;+\;\frac{\lambda}{2}\min_{g_{i}}\mathbb{E}\|g_{i}(Z_{i},C_{i})-X_{i}\|^{2}.$ (6)
@@
When $C_{1}$ is fixed, maximizing ${I(Z_{1},C_{1};X_{1})}=I(Z_{1};X_{1}\mid C_{1}){+I(X_{1};C_{1})}$ is equivalent to maximizing $I(Z_{1};X_{1}\mid C_{1})=H(X_{1}\mid C_{1})-H(X_{1}\mid Z_{1},C_{1})$, which is further equivalent to minimizing the conditional entropy $H(X_{1}\mid Z_{1},C_{1})$. By Fano’s inequality (Cover,, 1999, Theorem 8.6.6), the reconstruction error upper bounds this conditional entropy for any $g_{1}$:
@@ 3.3 Comparison with two SOTA Methods
Factorized Contrastive Learning (Factorized CL, Liang et al.,, 2023).
Like IndiSeek, this method uses the NCE-CLUB loss to upper bound the entanglement enforcing term. However, it replaces the complementary information capture term with an InfoNCE lower bound. For comparison, we consider its task-agnostic, two-step variant: [^5] Liang et al., (2021) also proposes a self-supervised variant of Factorized CL by leveraging task-relevant augmentations, more discussions on which are deferred to Section 6.
${\mathcal{L}}_{\rm FactorizedCL}(Z_{1};C_{1},\tau,\lambda)={\mathcal{L}}_{\rm NCE-CLUB}(Z_{1};C_{1})+\frac{\lambda}{2}{\mathcal{L}}_{\rm infoNCE}(h(C_{1},Z_{1}),X_{1},\tau).$ (7)
Disentangled Self-Supervised Learning (InfoDisen Wang et al., 2024a, ).
@@
${\mathcal{L}}_{\rm InfoDisen}(Z_{1};C_{1},\tau,\lambda)=\mathbb{E}_{(Z_{1},C_{1})}\left[\langle\mu(Z_{1}),\mu(C_{1})\rangle\right]+\frac{\lambda}{2}{\mathcal{L}}_{\rm infoNCE}(h(C_{1},Z_{1}),X_{1},\tau).$ (8)
Both methods inherit limitations: InfoNCE is sensitive to temperature tuning $\tau$ and may underperform as a surrogate for capturing complementary information, while the orthogonal loss assumes linearity and fails to capture nonlinear dependencies. IndiSeek avoids these issues by pairing NCE-CLUB with a reconstruction-based mutual information bound.
@@ 4 Experiments with a CITE-seq dataset
In this section, we evaluate the performance of IndiSeek on applications in single-cell biology. We consider the bone marrow CITE-seq dataset from Stuart et al., (2019) which consists of measurements in two modalities on $30672$ individual cells: transcriptome (RNA) and $25$ cell-surface proteins (ADT). We randomly split the dataset into a training set of size $15000$ and a test set of size $15672$. To leverage both RNA and ADT data for annotating cell types, Hao et al., (2021) proposed to cluster cells according to a weighted-nearest-neighbor (WNN) graph with cells as nodes. In its construction, the similarity between a cell $i$ and every other cell $j$ is measured by a weighted average of their respective similarities in the two modalities. The weights used by cell $i$ are in turn determined by the cross-modal predictive powers of its individual modalities after local smoothing. Thus, the RNA weight of each cell can be viewed as a quantification of the importance RNA plays in determining its cell state. The larger, the more important. Based on clustering nodes of the WNN graph, Hao et al., (2021) annotated cells at two granularity levels with $5$ (level-1) and $27$ (level-2) cell types, respectively.
@@
In the left panel of Figure 4, $27$ level-2 cell types are ranked according to average $\theta(c)$ scores of cells in the test set with respective cell type annotations, in which a higher rank indicates a higher impact of the RNA-specific information in determining the cell type. Each bar is colored according to the coarser level-1 cell type that each level-2 cell type belongs to for easier visual inspection. Different cell types within the T cell family have negative scores, indicating RNA-specific information is less informative compared to ADT-specific counterpart for differentiating different T cell subpopulations, which is in line with the observation that clustering transcriptomics data typically does not delineate T cell subpopulations (Szabo et al.,, 2019; Zheng et al.,, 2021) and gating based on surface protein markers is often needed (Kotliar et al.,, 2025). In contrast, RNA-specific information plays a more important role in distinguishing different progenitor cells than ADT-specific information. To further demonstrate the quality of learned modality-specific representations, we plot the ranks of average $\theta(c)$ values against the benchmark ranks of average RNA weights in Hao et al., (2021) of the 27 level-2 cell types in the right panel in Figure 4, which exhibits a strong correlation that is confirmed by Spearman’s rank correlation (Spearman,, 1961) of $0.922$. As a comparison, we repeated the foregoing experiments across $10$ random seeds while replacing IndiSeek with InfoDisen or FactorizedCL when training the same neural net architecture and the resulting alignment metrics with respect to the results in Hao et al., (2021) are presented in Figure 5. Both alternative methods yielded inferior alignment with the relative importance of the RNA modality in Hao et al., (2021) compared to IndiSeek. To ablate the choice of $\lambda$, in all experiments with real-world data, we report results with varying $\lambda$ on grid values $10^{j}$ for j=3,2,1,0,1,2,3j=-3,-2,-1,0,1,2,3.

### v3

@@ Abstract
Learning disentangled representations is a fundamental task in multi-modal learning. In modern applications such as single-cell multi-omics, both shared and modality-specific features are critical for characterizing cell states and supporting downstream analyses. Ideally, modality-specific features should be independent of shared ones while also capturing all complementary information within each modality. This tradeoff is naturally expressed through information-theoretic criteria, but mutual-information-based objectives are difficult to estimate reliably, and their variational surrogates often underperform in practice. In this paper, we introduce IndiSeek, a novel disentangled representation learning approach that addresses this challenge by combining an independence-enforcing objective with a computationally efficient reconstruction loss that bounds conditional mutual information. This formulation explicitly balances independence and completeness, enabling principled extraction of modality-specific features. We demonstrate the effectiveness of IndiSeek on synthetic simulations, a CITE-seq dataset and multiple real-world multi-modal benchmarks.
@@ 1 Introduction
The growing availability of multi-modal data has broadened the scope of representation learning. One notable example is the development of single-cell multi-omics technologies in genomics (teichmann2020method). For instance, CITE-seq (Cellular Indexing of Transcriptomes and Epitopes by Sequencing) (stoeckius2017simultaneous) enables simultaneous measurement of gene expression via single-cell RNA sequencing and protein abundance via antibody-derived tags (ADTs). Integrative analysis of both modalities at the single-cell resolution has improved tasks such as cell state delineation and enabled cell atlas construction and querying at finer granularity. Multi-modal data is also abundant at the intersection of computer vision, natural language processing, and audio processing, where combining information from different sources for representation learning forms the backbone of many groundbreaking progresses (radford2021learning; baltruvsaitis2018multimodal; jia2021scaling; akbari2021vatt; liang2024foundations).
@@
Shared features enhance predictive power across modalities by capturing cross-modal dependence. From an information-theoretic perspective, they should contain all information common across modalities. At the same time, preserving modality-specific diversity is equally important. In single-cell multi-omics, modality-specific features often correspond to unique biological signals that aid cell type identification and the discovery of novel or rare cell populations (hao2021integrated; caron2025multimodal). After all, if all the information in a certain modality can be inferred from other modalities, there is little reason to measure it in the first place. Finally, disentangling shared and modality-specific components improves interpretability: shared features summarize cross-modal information, while modality-specific features capture complementary signals.
Methodologically, shared information is usually extracted by maximizing cross-modal mutual information (tosh2021contrastive; sridharan2008information), with contrastive methods like CLIP (radford2021learning) widely adopted in practice. Under certain conditions, CLIP-learned representations have been shown to be not only sufficient (i.e., capturing all shared information), but also minimally sufficient (i.e., containing minimal extra information) (gui2025multi; oko2025statistical; lin2025statistical) under certain conditions.
In contrast, identifying modality-specific features, also known as disentangled representation learning, has become an active research area (fischer2020conditional; liang2023factorized; liu2023focal; wang2024information; dufumier2024align; wang2024decoupling). While shared features should be minimally sufficient, modality-specific features must both remain independent of (i.e., disentangled from) shared features and capture all complementary information within each modality (liang2023factorized; wang2024information; wang2024decoupling). From an information-theoretic perspective, this requires (1) simultaneously minimizing the mutual information between shared and modality-specific features and (2) maximizing the mutual information between data and their union. Multiple state-of-the-art methods (liang2023factorized; wang2024information) tackle this problem from such a perspective, but differ in how they approximate the underlying mutual information. Since such quantities are difficult to estimate with finite samples, these methods often trade off independence against completeness. As we show in the motivating examples below, this tradeoff limits their ability to achieve both goals simultaneously.
@@ 1.1 Motivating examples and limitations of SOTA
For any learned map $h_{1}$, we quantify the importance of each coordinate of $X_{1}$ in determining its output via gradients based on feature masking [^1] See Appendix A.1 for details on this feature importance metric. on an independent test set of size $1000$, following standard model-free importance measures (robnik2008explaining; zeiler2014visualizing; li2016understanding)
@@
We present the average coordinate importance metrics in learned modality-specific representation maps over $50$ simulation runs in Figures 1 and 2 for three different learning algorithms in the two settings, respectively. In both figures, the left panels report results of the SOTA approach FactorizedCL in liang2023factorized (without self-supervision), the middle panels report results of the SOTA approach InfoDisen in wang2024information, and the right panels report results of the new approach IndiSeek we propose in this paper. Tuning parameters [^2] See Appendix A.2 for an ablation study on performances with different tuning parameter values. are set at $0.01$ for InfoDisen and IndiSeek and at $1$ by default for FactorizedCL.
As we have reasoned when introducing the settings, the ideal modality-specific representations in Setting 1 should depend on and only on the last four coordinates, while in Setting 2, on and only the last two coordinates. The left and the middle panels of both figures suggest that neither FactorizedCL nor InfoDisen learn an ideal representation in either setting, as irrelevant coordinates play non-trivial roles in determining the learned modality-specific representations. Therefore, both SOTA approaches have their limitations in achieving the desired goal of disentangled representation learning. In contrast, the right panels of Figures 1 and 2 show that the new method IndiSeek we are to propose achieves it in both settings: the learned modality-specific features depend on and only on all coordinates that are independent of those involved in shared features.
@@ 3 Information-guided disentangled representation seeking (IndiSeek)
Given learned shared features, the desiderata in Section 2 motivate the search for disentangled modality-specific features, but leave open the question of how to realize them. Prior work has proposed an information-theoretic objective:
@@ 3.1 Step 1: Extracting Shared Features via CLIP
As noted at the end of Section 2, shared feature extraction is commonly handled by contrastive learning methods such as CLIP (radford2021learning), which we also adopt. Concretely, encoders $(f_{1},f_{2})$ are trained with the InfoNCE loss (oord2018representation), which provides a lower bound on the mutual information between $X_{1}$ and $X_{2}$. This encourages $C_{1}=f_{1}(X_{1})$ and $C_{2}=f_{2}(X_{2})$ to capture all cross-modal dependence while avoiding modality-specific redundancy.
@@ 3.2 Step 2: Extracting Modality-Specific Features
Given the shared features $(C_{1},C_{2})$, we aim to extract the modality-specific components $(Z_{1},Z_{2})$ by operationalizing the constrained problem in Eq. (3). Here, we note that, if the learned shared representations are redundant, that is, $C_{1}$ contains more information than the underlying shared features, learning the unique feature $Z_{1}$ by encouraging the independence between $Z_{1}$ and $C_{1}$ is lossy: the learned $Z_{1}$ may lose modality-specific information. To address this, we propose the cross-modal disentanglement:
$\min_{Z_{i}}\;\underbrace{I(Z_{i};C_{3-i})}_{\text{disentanglement-enforcing term}}-\quad\lambda\cdot\underbrace{I(Z_{i},C_{i};X_{i})}_{\text{complementary information capture term}}.$ (4)
The key is in the disentanglement term $I(Z_{1},C_{2})$ (or $I(Z_{2},C_{1})$). Even when the learned “shared ” information is redundant, i.e., when $C_{1}$ and $C_{2}$ both contain the underlying shared features (say, $C^{*}$), the redundancy is only a function of modality-specific features in $X_{2}$ that is independent of $X_{1}$; thus, given that $I(C^{*};Z_{1})=I(C_{2};Z_{1})$, we ought to encourage the independence between $Z_{1}$ and $C_{2}$ instead.
@@
  • For the disentanglement enforcing term $I(Z_{i};C_{3-i})$, which we aim to minimize, we replace it by an upper bound given by the NCE-CLUB loss (cheng2020club; liang2023factorized):
    ${\mathcal{L}}_{\rm NCE-CLUB}(Z_{1};C_{2})=\mathbb{E}_{(Z_{1},C_{2})}\left[\log p(Z_{1}\mid C_{2})\right]-\mathbb{E}_{Z_{1},\widetilde{C}_{2}}\left[\log p(Z_{1}\mid\widetilde{C}_{2})\right],$ (5)
    where $p(z_{1}\mid c_{2})$ is the underlying conditional density of $Z_{1}\mid C_{2}$ and $\widetilde{C}_{2}$ is an independent copy of $C_{2}$ from the same marginal distribution. [^4] In implementation, following cheng2020club, we replace conditional distributions by their estimates via multivariate Gaussian distributions whose moments are fitted from data in each epoch.
@@
$\mathcal{L}_{\text{{IndiSeek}}}(Z_{i};C_{1},C_{2},\lambda)\;=\;\mathcal{L}_{\text{NCE-CLUB}}(Z_{i};C_{3-i})\;+\;\frac{\lambda}{2\mathbb{E}\|X_{i}\|^{2}}\min_{g_{i}}\mathbb{E}\|g_{i}(Z_{i},C_{i})-X_{i}\|^{2}.$ (6)
@@
When $C_{1}$ is fixed, maximizing ${I(Z_{1},C_{1};X_{1})}=I(Z_{1};X_{1}\mid C_{1}){+I(X_{1};C_{1})}$ is equivalent to maximizing $I(Z_{1};X_{1}\mid C_{1})=H(X_{1}\mid C_{1})-H(X_{1}\mid Z_{1},C_{1})$, which is further equivalent to minimizing the conditional entropy $H(X_{1}\mid Z_{1},C_{1})$. By Fano’s inequality (cover1999elements, Theorem 8.6.6), the reconstruction error upper bounds this conditional entropy for any $g_{1}$:
@@ 3.3 Comparison with two SOTA Methods
Factorized Contrastive Learning (Factorized CL, liang2023factorized).
Like IndiSeek, this method uses the NCE-CLUB loss to upper bound the entanglement enforcing term. However, it replaces the complementary information capture term with an InfoNCE lower bound. For comparison, we consider its task-agnostic, two-step variant:
${\mathcal{L}}_{\rm FactorizedCL}(Z_{1};C_{1},C_{2},\tau,\lambda)={\mathcal{L}}_{\rm NCE-CLUB}(Z_{1};C_{1})+\frac{\lambda}{2}{\mathcal{L}}_{\rm infoNCE}(h(C_{1},Z_{1}),X_{1},\tau).$ (7)
liang2021multibench also proposes a self-supervised variant of Factorized CL by leveraging task-relevant augmentations, more discussions on which are deferred to Section 6 and Appendix A.1. In this paper, we adapt the unsupervised version of Factorized CL, i.e., we use CLUBInfoNCECritic between $Z_{1}$ and $C_{1}$ in liang2023factorizedgithub as the disentanglement objective, and adopt the implementation of CLUB loss in cheng2020club in both IndiSeek  and Factorized CL for fair comparison. (more details are presented in Appendix A.1).
Disentangled Self-Supervised Learning (InfoDisen wang2024information).
@@
${\mathcal{L}}_{\rm InfoDisen}(Z_{1};C_{1},C_{2},\tau,\lambda)=\mathbb{E}_{(Z_{1},C_{1})}\left[\langle\mu(Z_{1}),\mu(C_{1})\rangle\right]+\frac{\lambda}{2}{\mathcal{L}}_{\rm infoNCE}(h(C_{2},Z_{1}),X_{1},\tau).$ (8)
We note that, in the reconstruction term, wang2024information adopts the shared feature from the other modality, i.e., the actual reconstruction of interest becomes $I(Z_{i},C_{3-i};X_{i})$ instead of $I(Z_{i},C_{i};X_{i})$.
Both methods inherit limitations: InfoNCE is sensitive to temperature tuning $\tau$ and may underperform as a surrogate for capturing complementary information, while the orthogonal loss assumes linearity and fails to capture nonlinear dependencies. Additionally, without cross-modal disentanglement, InfoDisen  and FactorizedCL may fail to fully capture modality-specific features due to redundant shared features. IndiSeek avoids these issues by pairing NCE-CLUB with a reconstruction-based mutual information bound.
@@ 4 Experiments with a CITE-seq dataset
In this section, we evaluate the performance of IndiSeek on applications in single-cell biology. We consider the bone marrow CITE-seq dataset from stuart2019comprehensive which consists of measurements in two modalities on $30672$ individual cells: transcriptome (RNA) and $25$ cell-surface proteins (ADT). We randomly split the dataset into a training set of size $15000$ and a test set of size $15672$. To leverage both RNA and ADT data for annotating cell types, hao2021integrated proposed to cluster cells according to a weighted-nearest-neighbor (WNN) graph with cells as nodes. In its construction, the similarity between a cell $i$ and every other cell $j$ is measured by a weighted average of their respective similarities in the two modalities. The weights used by cell $i$ are in turn determined by the cross-modal predictive powers of its individual modalities after local smoothing. Thus, the RNA weight of each cell can be viewed as a quantification of the importance RNA plays in determining its cell state. The larger, the more important. Based on clustering nodes of the WNN graph, hao2021integrated annotated cells at two granularity levels with $5$ (level-1) and $27$ (level-2) cell types, respectively.
@@
In the left panel of Figure 4, $27$ level-2 cell types are ranked according to average $\theta(c)$ scores of cells in the test set with respective cell type annotations, in which a higher rank indicates a higher impact of the RNA-specific information in determining the cell type. Each bar is colored according to the coarser level-1 cell type that each level-2 cell type belongs to for easier visual inspection. Different cell types within the T cell family have negative scores, indicating RNA-specific information is less informative compared to ADT-specific counterpart for differentiating different T cell subpopulations, which is in line with the observation that clustering transcriptomics data typically does not delineate T cell subpopulations (szabo2019single; zheng2021pan) and gating based on surface protein markers is often needed (kotliar2025reproducible). In contrast, RNA-specific information plays a more important role in distinguishing different progenitor cells than ADT-specific information. To further demonstrate the quality of learned modality-specific representations, we plot the ranks of average $\theta(c)$ values against the benchmark ranks of average RNA weights in hao2021integrated of the 27 level-2 cell types in the right panel in Figure 4, which exhibit a strong correlation that is confirmed by Spearman’s rank correlation (spearman1961proof) of $0.910$. In Appendix A.3, we also present the UMAP of learned representations, which exhibits clear clusters with respect to two levels of cell types. As a comparison, we repeated the foregoing experiments across $10$ random seeds while replacing IndiSeek with InfoDisen or FactorizedCL when training the same neural net architecture and the resulting alignment metrics with respect to the results in hao2021integrated are presented in Figure 5, in which we also report the error bars for results over $10$ random seeds. Both alternative methods yielded inferior alignment with the relative importance of the RNA modality in hao2021integrated compared to IndiSeek. To ablate the choice of $\lambda$, in all experiments with real-world data, we report results with varying $\lambda$ on grid values $10^{j}$ for j=3,2,1,0,1,2,3j=-3,-2,-1,0,1,2,3.